# Hierarchical Mixture of Experts: Generalizable Learning for High-Level Synthesis

### v3

@@ Introduction
With the heated demand for domain-specific accelerators, field-programmable gate arrays (FPGA) are widely used. It is, however, labor-intensive to write register-transfer-level hardware description languages, such as VHDL and Verilog. High-level synthesis (HLS) provides a much easier solution by compiling a source code written in a software programming language, such as C and MatLab, into an FPGA circuit (Cong et al. 2011, 2022; Schafer and Wang 2019).
@@
The source code consists of a program (also called “kernel”) that describes the FPGA’s functional behaviors, and several pragmas inserted in the program that instruct the hardware synthesis process, such as parallelization, pipeline, etc., as illustrated in the code snippets in Fig. 1. While it is relatively easy to design the program, it heavily requires hardware knowledge to design the pragmas, and different pragma designs can lead to significantly different FPGA performances, posing a great challenge for software developers (Sohrabizadeh et al. 2022). Recent works have automated pragma design for programs, employing heuristics methods or machine learning methods. The heuristic methods use bottleneck analysis (Sohrabizadeh et al. 2021) or non-linear programming based on lower-bound objective function (Pouget, Pouchet, and Cong 2024a, b). The machine learning methods train a surrogate model to predict the FPGA’s performance from the source code, so that we can rely on the model prediction to find the best pragmas without running time-consuming HLS. Since HLS data is very scarce, it is difficult to train a large language model, and graph neural networks (GNNs) that are based on the control data flow graph (Sohrabizadeh et al. 2022; Bai et al. 2022; Sohrabizadeh et al. 2023; Ustun et al. 2020; Wu, Xie, and Hao 2023; Wu et al. 2022a) are widely utilized. Some very recent work (Qin et al. 2024) explores using both GNN and lightweight language models.
@@
While machine learning models have better learning ability than the heuristic methods, they do not generalize well to unseen kernels. In real-world applications, we often encounter new circuit design requirements. The model trained on existing kernels usually fails to perform well on new kernels. This can be viewed as a domain generalization (Bai et al. 2022; Kwon and Carloni 2020) problem, where each kernel is a domain. Note it is very time-consuming to run HLS to acquire the labels on new kernels, with each run taking minutes to hours (Sohrabizadeh et al. 2023), thus data-efficient fine-tuning is required. Domain generalization methods for GNN usually employ adversarial training to align the representation space between different domains (Dai et al. 2019; Zhang et al. 2019; Wu et al. 2020; Shen et al. 2023), or do data augmentation to learn invariant representation under risk extrapolation (Wu et al. 2022b; Liu et al. 2023a). However, these approaches are not applicable to the HLS prediction task, as the differences between kernels are informative for the prediction, so forcing them to the same distribution does not work.
@@
Nonetheless, a unique opportunity of the HLS prediction task is that kernels usually share some similar substructures, as shown in Fig. 1. It could be beneficial if we can leverage such similarities. We thus propose a two-level hierarchical mixture of experts (MoE) to address this issue. MoE (Jacobs et al. 1991; Shazeer et al. 2017) uses several expert networks instead of a single fixed network, and it computes the weighted sum of their outputs. Different experts can specialize in different regions in the hidden representation space, so that similar substructures in different domains can be processed by the same experts. A gating network computes the expert assignment, which learns which expert should deal with which substructure. GraphMETRO (Wu et al. 2024) applies MoE to improve GNNs’ generalizability by predefining a set of domain shifts and training each expert to deal with one type of domain shift by data augmentation. We take an orthogonal direction, eliminate the need for hardware knowledge to define the domain shifts, and focus on the model structure rather than data augmentation.
@@
Another unique opportunity for HLS prediction is that the input graph naturally has three levels of granularity: nodes, blocks, and the whole graph. MoE can be constructed on either of these three granularities. A natural question is which granularity to use. It is a hard task as different kernels benefit from different granularities. Simply applying MoE on all three granularities in one model does not work. Therefore, we propose a higher-level MoE to aggregate the three low-level MoEs according to the need. Nonetheless, optimization with the hierarchical MoE is challenging. Since the three low-level MoEs are different, training easily leads to expert polarization. We propose a two-stage training to stabilize the training. To the best of our knowledge, while some papers (Shazeer et al. 2017) have studied hierarchical MoE to improve memory efficiency, this is the first study that demonstrates the performance gain of hierarchical MoE. We conduct extensive experiments on the largest HLS benchmark dataset, HLSyn (Bai et al. 2023), and the experiment results reveal the effectiveness of hierarchical MoE. In summary, we make the following contributions:
@@
  • Extensive experiments verify the effectiveness of our methods, where the average speedup of the FPGA design is 26.6% higher than that of the baseline method.
@@
There are two directions for HLS pragma design automation. The first direction is heuristic-driven. AutoDSE (Sohrabizadeh et al. 2021) uses bottleneck-guided searching based on the HLS feedback, and it is well-generalizable. Non-linear programming performs well on affine kernels based on a lower-bound objective function (Pouget, Pouchet, and Cong 2024a, b). The second direction is data-driven, which trains a surrogate machine learning model to predict the FPGA’s latency and resource utilization. GNNs (Sohrabizadeh et al. 2022, 2023; Ustun et al. 2020; Wu et al. 2022a; Wu, Xie, and Hao 2023, 2021; Qin et al. 2024) are widely used. While machine learning models could perform better than heuristic methods with sufficient training data, they suffer from poor generalizability. Several previous works have tried to solve this issue. (Ferretti et al. 2020) transforms the best pragma design from the most similar source kernel to the target kernel, but its transformation is too simple and highly limited. (Kwon and Carloni 2020) trains a separate domain adaptor for each domain, while (Bai et al. 2022) uses the meta-learning method, MAML (Finn, Abbeel, and Levine 2017), to find a generalizable parameter initialization. However, since we have many kernels, it is difficult to train MAML to get stable results. We only have hundreds of data for each kernel, so it is difficult to train an adaptor. Instead, we explore an orthogonal direction: to improve the model structure’s generalizability.
@@
HLS prediction is formalized as a graph regression task. Following previous works (Sohrabizadeh et al. 2022, 2023), we utilize ProGraML (Cummins et al. 2021) graph to represent a source code, which is a control data flow graph. Nodes represent instructions, variables, and constant values, and edges represent the control flow and data flow. A GNN model is trained to predict the FPGA’s latency and the utilization of four resources: LUT, FF, DSP, and BRAM. In the domain generalization setting, we train the model on $N$ source kernels $D^{(train)}=\{D_{1},D_{2},...,D_{N}\}$, where $D_{i}=\{(X_{i},T_{i1},Y_{i1}),(X_{i},T_{i2},Y_{i2}),...,(X_{i},T_{ii_{n}},Y_{ii_{n}})\}$ is the $i$-th kernel containing $i_{n}$ samples, and each sample consists of a program $X$, a pragma design $T$, and a label $Y$. For a new kernel $D_{test}$, we consider the constraint of data scarcity where we only use $k$ samples from the dataset $D^{(k)}\subset D_{test},|D^{(k)}|=k$, to fine-tune the model. After fine-tuning, there are two ways of evaluation: offline evaluation and online evaluation. In the offline evaluation, we evaluate the MSE on the left-out test samples: $D_{test}\setminus D^{(k)}$. In the online evaluation, we do design space exploration (DSE) to search for good pragma designs. Based on the fine-tuned model’s prediction, we select the top $M$ designs to run HLS, and evaluate the speedup ($M$ is set to 10 in our experiments).
@@
The proposed hierarchical MoE can be adapted to any GNN model. We use one of the SOTA GNN models for HLS prediction, HARP (Sohrabizadeh et al. 2023), as the base model. Since the original ProGraML graph is not good at modeling long-range dependency, HARP creates a “pseudo node” for each basic block and connects it with every node within that basic block. A basic block is a sequence of instructions with a single entry point and a single exit point where the terminator instruction can be a branch, return, etc. We illustrate HARP’s model structure in the appendix. We use $V$ to denote the set of all nodes and $V_{B}$ to denote the set of pseudo nodes. HARP consists of five components: (1) GNN encoder: it consists of 6 GNN layers and learns node representation; (2) Pragma MLP: it uses an MLP for each pragma type to transform the representation of pseudo nodes that are modified by that pragma type; (3) Another GNN layer after pragma MLP: it further updates the node representations; (4) Graph pooling: it performs graph pooling on pseudo nodes $V_{B}$ to form one graph representation and on normal nodes $V\setminus V_{B}$ to form another, then concatenates the two; (5) Output MLP: it makes the final prediction.
@@ Methods
A unique opportunity for the HLS prediction task is that the input graph has three granularities: normal nodes that represent data and instructions, pseudo nodes that represent basic blocks, and the whole graph that represents a source code file. If a data point from the target kernel shares a similar statement (node), basic block, or the whole code (graph) with a data point from the training kernels, MoE could be helpful. Thus, we consider employing MoE on the three granularities: MoE on all nodes including normal nodes and pseudo node (node MoE), MoE only on pseudo nodes (block MoE), and MoE on the graph (graph MoE). Based on our pre-exploration, the best practice is to apply MoE in the components after the pragma MLP, since the experts need to share the same encoder and pragma MLPs. In our pre-exploration, we found that the best structure of the gating network is a linear layer. The model structures are shown in Figure 2.
@@
where $\boldsymbol{W_{G}^{node}}\in\mathbb{R}^{n\times d}$ is the parameter of the gating network, $N(v_{i})$ is the neighbors of $v_{i}$, and $GNN_{j}(\cdot)$ is the $j$-th expert. The $softmax$ function is a commonly-used normalization method that ensures the weights of all experts sum up to 1. The expert weights can be seen as “attention”. We calculate the expert weights (attention scores) using the gating network based on the node embeddings $\boldsymbol{h_{i}}$. By employing the MoE layer, different nodes can utilize different parameters for message passing and message aggregation.
Block MoE. We have tried applying MoE on the pragma MLP, but the performance has decreased. It might be because each pragma type has a separate pragma MLP, so the training data for each pragma MLP is already limited. If we apply MoE, it will dilute the training data even more. Instead, we add an additional MoE layer on the pseudo nodes’ representations after the pragma MLP, so that we can increase the expressiveness of the pseudo nodes’ representations before graph pooling. The representation space of the pseudo nodes after the block MoE could be different from normal nodes, so it is no longer suitable to have another GNN layer after the pragma MLP to do message-passing between pseudo nodes and normal nodes. We have also tried applying an additional layer for normal nodes after the pragma MLP, but the performance is unsatisfactory. It might be because the normal nodes lack a global view, so it is not useful to further transform its representation. After having the block MoE, performing graph pooling only on the pseudo nodes performs better than graph pooling on both pseudo nodes and normal nodes. Therefore, we do not utilize normal nodes after the GNN encoder in the block MoE model. Ablation study of various designs is in the appendix.
@@
Graph MoE. We apply the graph MoE on the output MLP. We denote the graph representation after graph pooling as $\boldsymbol{h_{G}}$. Different from the block MoE structure, here we use the original pragma MLP and graph pooling structures as HARP. We employ the output MLP as the expert network. The graph MoE is formulated as:
$\hat{Y}^{(t)}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}^{graph}}\boldsymbol{h_{G}})_{j}\cdot MLP^{(t)}_{j}(\boldsymbol{h_{G}}),$ (3)
where $W_{G}^{graph}\in\mathbb{R}^{n\times d}$ is the parameter of the gating network, $MLP^{(t)}_{j}(\cdot)$ is the $j$-th expert for predicting the objective $t$ ($t$ could be latency or a certain resource’s utilization). The expert assignment is the same for all prediction objectives. The MLP contains four linear layers and the ELU activation function in between. Different design points can utilize different experts to make the final prediction.
Regularization term. As discovered in many literature (Shazeer et al. 2017; Wang et al. 2023; Li et al. 2023), expert polarization is a common problem of MoE. Due to the random initialization of expert networks, different experts initially perform differently. The gating network learns to assign higher weights to better experts. It results in more training of the initial better experts, leading to their even better performance, which forms a cycle. As training continues, the MoE model might collapse and only use the best expert. To avoid this issue, we apply a regularization term commonly used in MoE (Shazeer et al. 2017):
@@
We propose two designs of the high-level gating network, as illustrated in Figure 2. The first design is to perform graph pooling on the input node features to form a graph representation as the input to the high-level gating network. We use the self-attention graph pooling. Denoting the input feature of node $v_{i}$ as $\boldsymbol{x_{i}}$, the graph pooling is formalized as:
$\boldsymbol{x_{G}}=\sum_{i\in V}softmax(MLP(\boldsymbol{x_{i}}))\cdot\boldsymbol{x_{i}}$ (5)
where $X_{G}$ is the aggregated input feature, and $V$ is the set of all nodes. The second design is to concatenate the graph representation in the three low-level MoE models. The second design performs better in our experiments, since the hidden representations are more expressive than the input features. Nonetheless, when we use a sparse MoE where only one or two experts are selected, the first design will be more memory efficient. It can determine the expert assignment before the computation of three expert models, thus reducing unnecessary computation. However, the best-performing method is to utilize all the experts, and in this case, the two designs are similar in efficiency. We utilize all the experts in our main experiments. We also apply the regularization term for the high-level gating network.
@@
Table 2: Dataset statistics. “#Class” and “#Regre” denote the number of classification and regression data. The kernel acronyms represent “fdtd-2d-large”, “gemver-medium”, “syr2k”, “gemm-p”, “jacobi-2d”, and “trmm-opt”.
Models. Based on our pre-explorations, we use 4 experts in the low-level MoEs, and we set the regularization terms’ weights of both low-level and high-level MoEs to 5e-3. Our baselines include the SOTA GNN model, HARP (Sohrabizadeh et al. 2023); HARP+MAML (Bai et al. 2022), which applies meta-learning method MAML (Finn, Abbeel, and Levine 2017) on HARP inspired by (Bai et al. 2022) to learn a more generalizable initialization of parameters; and ProgSG (Qin et al. 2024), which combines HARP and language model. Although hierarchical MoE brings many more parameters, it needs about 32 GB of memory and can still fit into a single GPU. HARP and our model contain 359,370 and 1,329,403 parameters respectively.
Evaluation. We train the regression and classification models on the source kernels. We want to mimic the domain transfer situation of having scarce but representative labeled data on the target kernels, so we use 50 data points per kernel to fine-tune our regression model, and roughly the same ratio of data points, 265 samples per kernel, to fine-tune the classification model. To select representative data points, we use K-means based on the graph representation. We conduct both offline and online evaluations. In offline evaluation, we calculate the fine-tuned regression model’s mean squared error (MSE) on the left-out data points in the target kernels for each regression objective and sum up the five objectives’ MSE. In the online evaluation, we use the DFS search used in previous studies (Sohrabizadeh et al. 2022, 2023) to search for pragma designs. We limit our search range to 75,000 pragma designs, since it typically takes an hour. We use the fine-tuned classification model to predict the validity and the fine-tuned regression model to predict the latency and resource utilization. From designs that are predicted to be valid and satisfy resource constraints, we choose the top 10 designs with the lowest predicted latencies to run HLS. We report the best design from the selected top-10 and the training dataset of the target kernels during fine-tuning, since these are all labeled data points, and we calculate its speedup compared to the best design in the dataset. We run the offline evaluation five times and the online evaluation three times, then we report the mean results.
@@
Table 1 shows the main results. In the main experiments, we use the second design of the high-level gating network as it performs better. MAML could not perform well in this setting. It might be because we have many more source kernels compared to the paper that proposes to use MAML on this task (Bai et al. 2022). Different kernels might result in different directions of the meta gradient, leading to unstable training. ProgSG which combines GNN and language models has a strong ability when the training data is sufficient, but it overfits in the data-scarce setting. During fine-tuning, the training loss is lower than 0.05, but the test loss is high. Comparatively, the hierarchical MoE is more generalizable. In the online evaluation, on most kernels, the best of three single MoE models outperforms HARP. However, different kernels favor different low-level MoEs. For example, the block is the best granularity for “Fd”, while the graph is the best granularity for “Gemv”, “Ja”, and “Tr”. By aggregating them together, hierarchical MoE performs the best or close to the best on almost every kernel.
@@
Complex kernels. Kernels in the HLSyn dataset are generally small. We further test the models on 5 more complex kernels, including ”3d-rendering” and ”spam-filter” from the Rosetta benchmark (Zhou et al. 2018) and three self-constructed kernels, which are introduced in the appendix. They are closer to real-world applications. Following previous settings, we train the regression model on all HLSyn kernels and finetune it on 50 points per kernel on the new kernels. The results are shown in Table 3. It is challenging for all models to perform well, while the hierarchical MoE model performs the best. It remains a future work to generally enhance model capacity on complex kernels.
@@
Table 4: Using MoE on various granularities in a single model. “N”, “B”, and “G” represent node, block, and graph.
To verify the effectiveness of our model design, we run extensive experiments on various model structures as ablation studies. First, instead of aggregating the three low-level MoEs, can we apply MoE on the three granularities together in a single model? We specify the detailed structure design in the appendix. Table 4 shows the results. When we use MoE on two or three granularities in a single model, the loss is usually higher than the lowest loss when only using MoE on one granularity. Apart from selecting 50 representative data points by K-means for fine-tuning, we also experiment with random data split. We use the same random split for all models to ensure fair comparison. Hierarchical MoE is the most stable model when it faces low-quality fine-tuning data in the random split. If we stack MoE on various granularities in the same model, the structure might be too complex and thus unstable to train on a small dataset.
@@
Second, does the hierarchical MoE benefit from combining the three granularities? To answer it, we still use the hierarchical MoE structure, but we use the same low-level MoE model as the three experts of the high-level MoE. Table 5 shows the results. It verifies that aggregating the three granularities performs the best. Using only the node MoE or graph MoE also performs well in the online evaluation, and this might be due to the increased number of experts. Nonetheless, aggregating the three granularities could further improve the performance.

### v4

@@ Introduction
With the heated demand for domain-specific accelerators, field-programmable gate arrays (FPGAs) are widely used. It is, however, labor-intensive to write register-transfer-level hardware description languages, such as VHDL and Verilog. High-level synthesis (HLS) provides a much easier solution by compiling a source code written in C++ into an FPGA circuit (Cong et al. 2011, 2022; Schafer and Wang 2019).
@@
The source code consists of a C++ program (also called a “kernel”) that describes the FPGA’s functional behaviors and several pragmas inserted into the program. The pragmas instruct the hardware synthesis process, such as parallelization, pipeline, and loop tiling, as illustrated in the code snippets in Fig. 1. Different pragmas can lead to significantly different FPGA performances, so it is very important fo find the most effective pragmas. It still requires a great deal of hardware knowledge to design the pragmas, posing a great challenge for software developers (Sohrabizadeh et al. 2022).
@@
Recent works have designed several methods to automate pragma design for a given C++ kernel. Heuristic methods use bottleneck analysis (Sohrabizadeh et al. 2021) or non-linear programming based on lower-bound estimation (Pouget, Pouchet, and Cong 2024a, b). Machine learning methods train a surrogate model to predict the FPGA’s performance from the source code, so that we can rely on the model prediction to find the best pragmas without running time-consuming HLS. HLS data is scarce, so it is difficult to train a large language model, and graph neural networks (GNNs) based on the control data flow graph are widely used (Sohrabizadeh et al. 2022; Bai et al. 2022; Sohrabizadeh et al. 2023; Ustun et al. 2020; Wu, Xie, and Hao 2023; Wu et al. 2022a). A recent work (Qin et al. 2024) also uses both GNN and lightweight language models.
@@
While machine learning models have better learning abilities than heuristic methods, they suffer from poor generalizability. A trained model usually fails to perform well on new kernels. This can be viewed as a domain generalization problem (Bai et al. 2022; Kwon and Carloni 2020), where each kernel is one domain. The first challenge is that the domain difference is too big, and the difference is informative for the prediction. Previous domain generalization methods for GNN usually employ adversarial training to align the representation space between different domains (Dai et al. 2019; Zhang et al. 2019; Wu et al. 2020; Shen et al. 2023), or do data augmentation to learn invariant representation under risk extrapolation (Wu et al. 2022b; Liu et al. 2023a). They are not directly applicable to our task, since we do not want to force different domains’ representations to fit in the same distribution. The second challenge is that it is time-consuming to run HLS to acquire labels on new kernels, with each run taking minutes to hours (Sohrabizadeh et al. 2023).
@@
While the HLS prediction task is challenging to domain generalization, it also provides a unique opportunity: different kernels usually share some similar substructures. As shown in Fig 1, it would be beneficial if we could utilize the knowledge learned from similar substructures on the old kernels to apply on the new kernels. MoE (Jacobs et al. 1991; Shazeer et al. 2017) is an appropriate solution. It computes the weighted sum of the outputs from several expert networks instead of a single network. Different experts can specialize in different regions in the hidden representation space, so similar substructures in different domains can be processed by similar experts. GraphMETRO (Wu et al. 2024) has demonstrated MoE’s effect on improving GNNs’ generalizability. It defines several types of domain shifts and trains each expert to deal with one type by data augmentation. We take an orthogonal direction, eliminate the need to define the domain shifts, and focus on the model structure rather than data augmentation.
@@
The HLS graph naturally has three levels of granularity: nodes, blocks, and the whole graph. To build an MoE model suitable for this specific task, we consider constructing MoE on these three granularities. We find that different kernels benefit from applying MoE on different granularities. A natural question is which granularity to use. Simply applying MoE on all three granularities simultaneously does not work well. Therefore, we propose a higher-level MoE to aggregate the three low-level MoEs according to the need, which forms a two-level hierarchical MoE structure. It can flexibly adjust the weight of each low-level MoE based on the data. To the best of our knowledge, while some papers (Shazeer et al. 2017) have studied hierarchical MoE to improve memory efficiency, this is the first study to demonstrate the performance gain of hierarchical MoE.
Nonetheless, optimization with the hierarchical MoE is challenging. Different from previous MoE papers where all the expert networks have the same or similar structures, our three low-level MoEs have very different structures, and thus very different convergence speeds. As a result, training easily leads to expert polarization and mode collapse. It leads to only using the single expert that converges the fastest. Thus, we propose a two-stage training strategy to stabilize the training. We conduct extensive experiments on the benchmark dataset, HLSyn (Bai et al. 2023), and the experiment results reveal the effectiveness of hierarchical MoE. In summary, we make the following contributions:
@@
  • Extensive experiments verify our methods’ effectiveness, where the geometric mean speedup of the FPGA design is 26.6% higher than that of the baseline method.
@@
There are two directions for HLS pragma design automation. The first direction is heuristic-driven. AutoDSE (Sohrabizadeh et al. 2021) uses bottleneck-guided searching based on the HLS feedback, and it is well-generalizable. Another heuristic method is non-linear programming, which performs well on affine kernels based on a lower-bound objective function (Pouget, Pouchet, and Cong 2024a, b). The second direction is data-driven, which trains a surrogate machine learning model to predict the FPGA’s latency and resource utilization. GNNs are widely used (Sohrabizadeh et al. 2022, 2023; Ustun et al. 2020; Wu et al. 2022a; Wu, Xie, and Hao 2023, 2021; Qin et al. 2024).
While machine learning models perform better than heuristic methods with sufficient training data, they suffer from poor generalizability. Some previous works have tried to solve this issue. (Ferretti et al. 2020) transforms the best pragma design from the most similar source kernel to the target kernel, but its transformation is too simple and limited. (Kwon and Carloni 2020) trains a separate adaptor for each domain, while (Bai et al. 2022) uses the meta-learning method, MAML (Finn, Abbeel, and Levine 2017), to find generalizable parameter initializations. However, we have many kernels, so it is difficult to train MAML to get stable results. We only have hundreds of data points per kernel, so it is difficult to train an adaptor. We explore an orthogonal direction: to improve the model structure’s generalizability.
@@
HLS prediction is formalized as a graph regression task. Following previous works (Sohrabizadeh et al. 2022, 2023), we utilize the ProGraML graph to represent a source code (Cummins et al. 2021), which is a control data flow graph. Nodes represent instructions, variables, and constant values, and edges represent the control flow and data flow. A GNN model is trained to predict the FPGA’s latency and the utilization of four resources: LUT, FF, DSP, and BRAM. In the domain generalization setting, we train the model on $N$ source kernels $D^{(train)}=\{D_{1},D_{2},...,D_{N}\}$, where $D_{i}=\{(X_{i},T_{i1},Y_{i1}),(X_{i},T_{i2},Y_{i2}),...,(X_{i},T_{ii_{n}},Y_{ii_{n}})\}$ is the $i$-th kernel containing $i_{n}$ samples, and each sample consists of a program $X$, a pragma design $T$, and a label $Y$. For a new kernel $D_{test}$, we consider the constraint of data scarcity where we only use $k$ samples from the dataset $D^{(k)}\subset D_{test},|D^{(k)}|=k$, to fine-tune the model. After fine-tuning, there are two ways of evaluation: offline evaluation and online evaluation. In the offline evaluation, we evaluate the MSE on the left-out test samples: $D_{test}\setminus D^{(k)}$. In the online evaluation, we do design space exploration (DSE) to search for good pragma designs. Based on the fine-tuned model’s prediction, we select the top $M$ designs to run HLS, and evaluate the speedup ($M$ is set to 10 in our experiments).
@@
The proposed hierarchical MoE can be adapted to any GNN model. We use one of the SOTA GNN models for HLS prediction, HARP (Sohrabizadeh et al. 2023), as the base model. Since the original ProGraML graph is not good at modeling long-range dependency, HARP creates a “pseudo node” for each basic block and connects it with every node within that basic block. A basic block is a sequence of instructions with a single entry point and a single exit point where the terminator instruction can be a branch, return, etc. We illustrate HARP’s model structure in the appendix. We use $V$ to denote the set of all nodes and $V_{B}$ to denote the set of pseudo nodes. HARP consists of five components: (1) A GNN encoder that consists of 6 GNN layers and learns node representation; (2) pragma MLPs that use an MLP for each pragma type to transform the representation of pseudo nodes that are modified by that pragma type; (3) another GNN layer after pragma MLP that further updates the node representations; (4) graph pooling, which performs graph pooling on pseudo nodes $V_{B}$ to form one graph representation and on normal nodes $V\setminus V_{B}$ to form another, then concatenates the two; and (5) output MLP that makes the final prediction.
@@ Methods
A unique opportunity for the HLS prediction task is that the input graph has three granularities: normal nodes that represent data and instructions, pseudo nodes that represent basic blocks, and the whole graph that represents a source code file. When a data point from the target kernel shares a similar statement (node), basic block, or the whole code (graph) with a data point from the training kernels, then MoE is helpful. Thus, we consider employing MoE on the three granularities: MoE on all nodes including normal nodes and pseudo node (node MoE), MoE only on pseudo nodes (block MoE), and MoE on the graph (graph MoE). Based on our pre-exploration, the best practice is to apply MoE in the model components after the pragma MLP (the second component in HARP), since the experts need to share the same encoder and pragma MLPs. In our pre-exploration, we found that the best structure of the gating network is a linear layer. The model is shown in Figure 2.
@@
where $\boldsymbol{W_{G}^{node}}\in\mathbb{R}^{n\times d}$ is the parameter of the gating network, $N(v_{i})$ is the set of neighbors of $v_{i}$, and $GNN_{j}(\cdot)$ is the $j$-th expert. The $softmax$ function is a normalization method that ensures that the weights of all the experts sum up to 1. The expert weights can be seen as “attention.” We calculate the expert weights (attention scores) using the gating network based on the node embeddings $\boldsymbol{h_{i}}$. By employing the MoE layer, different nodes can utilize different parameters for message passing and message aggregation.
Block MoE. We have tried applying MoE on the pragma MLP, but the performance decreased. It might be because each pragma type has a separate pragma MLP, so the training data for each pragma MLP is already limited. After we applied MoE, it diluted the training data even more. Instead, we add an additional MoE layer on the pseudo nodes’ representations after the pragma MLP, so that we can increase the expressiveness of the pseudo nodes’ representations before the graph pooling. The representation space of the pseudo nodes after the block MoE might be different from normal nodes, so it may not be suitable to have another GNN layer after the block MoE to do message-passing between pseudo nodes and normal nodes. We have also tried applying an additional layer for normal nodes after the pragma MLP, but the performance is unsatisfactory. It might be because the normal nodes lack a global view, so it is not useful to further transform its representation. After having the block MoE, doing graph pooling only on the pseudo nodes performs better than doing graph pooling on both pseudo nodes and normal nodes. Therefore, we do not utilize normal nodes after the GNN encoder in the block MoE model. Ablation study of various designs is in the appendix.
@@
Graph MoE. We apply the graph MoE on the output MLP. We denote the graph representation after graph pooling as $\boldsymbol{h_{G}}$. Different from the block MoE structure, here we use the original model components of the HARP model except for the output MLP. We employ $n$ output MLPs as $n$ expert networks. The graph MoE is formulated as:
$\hat{Y}^{(t)}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}^{graph}}\boldsymbol{h_{G}})_{j}\cdot MLP^{(t)}_{j}(\boldsymbol{h_{G}}),\vspace{-1mm}$ (3)
where $W_{G}^{graph}\in\mathbb{R}^{n\times d}$ is the parameter of the gating network, and $MLP^{(t)}_{j}(\cdot)$ is the $j$-th expert for predicting the objective $t$ ($t$ could be latency or a certain resource’s utilization). The expert assignment is the same for all prediction objectives. The MLP contains four linear layers and the ELU activation function in between each. Different design points can utilize different experts to make the final prediction.
Regularization term. As discovered in previous research (Shazeer et al. 2017; Wang et al. 2023; Li et al. 2023), expert polarization is a common problem of MoE. Due to the random initialization of expert networks, different experts initially perform differently. The gating network learns to assign higher weights to better experts. It results in more training of the initial better experts, leading to their even better performance, which is self-reinforcing. As training continues, the MoE model might collapse and only use the best expert. To avoid this problem, we apply a regularization term commonly used in MoE (Shazeer et al. 2017):
@@
We tested two design choices of the high-level gating network, as illustrated in Figure 2. The first design is to perform graph pooling on the input node features to form a graph representation as the input to the high-level gating network. We use the self-attention graph pooling. Denoting the input feature of node $v_{i}$ as $\boldsymbol{x_{i}}$, the graph pooling is formalized as:
$\boldsymbol{x_{G}}=\sum_{i\in V}softmax(MLP(\boldsymbol{x_{i}}))\cdot\boldsymbol{x_{i}},$ (5)
where $X_{G}$ is the aggregated input feature, and $V$ is the set of all nodes. The second design is to concatenate the graph representation in the three low-level MoE models.
If we use a sparse MoE where only one or two experts are selected, the first design would be more memory efficient. The first design can determine the expert assignments before the computation of three expert models, thus reducing unnecessary computation. However, the best-performing method is to utilize all the experts, and in this case, the two designs are similar in efficiency. The second design performs better in our experiments, since the hidden representations are more expressive than the input features, and so we utilize the second design in our experiments. We also apply the regularization term for the high-level gating network.
@@
Table 2: Dataset statistics. “#Class” and “#Regre” denote the number of classification and regression data. The kernel acronyms represent “fdtd-2d-large,” “gemver-medium,” “syr2k,” “gemm-p,” “jacobi-2d,” and “trmm-opt.”
Models. Based on our pre-explorations, we use 4 experts in the low-level MoEs, and we set the regularization terms’ weights of both low-level and high-level MoEs to 5e-3. Our baselines include (1) the SOTA GNN model, HARP (Sohrabizadeh et al. 2023); (2) HARP+MAML (Bai et al. 2022), which applies the meta-learning method, MAML (Finn, Abbeel, and Levine 2017), on HARP inspired by (Bai et al. 2022) to learn a more generalizable initialization of parameters; and (3) ProgSG (Qin et al. 2024), which combines HARP and the language model. Hierarchical MoE brings many more parameters and needs about 32 GB of memory, but it can still fit into a single GPU. HARP and our model contain 359,370 and 1,329,403 parameters, respectively.
Evaluation. We train the regression and classification models on the source kernels. We want to mimic the domain transfer situation of having scarce but representative labeled data on the target kernels, so we use 50 data points per kernel to fine-tune our regression model, and roughly the same ratio of data points, 265 samples per kernel, to fine-tune the classification model. To select representative data points, we use K-means based on the graph representation. We conduct both offline and online evaluations. In offline evaluation, we calculate the fine-tuned regression model’s mean squared error (MSE) on the left-out data points in the target kernels for each regression objective and sum up the five objectives’ MSE. In the online evaluation, we use the DFS search used in previous studies (Sohrabizadeh et al. 2022, 2023) to search for pragma designs. We limit our search range to 75,000 pragma designs, since it typically takes an hour. We use the fine-tuned classification model to predict the validity, and we use the fine-tuned regression model to predict the latency and resource utilization. From designs that are predicted to be valid and satisfy resource constraints, we choose the top 10 designs with the lowest predicted latencies to run HLS. We report the best design from the selected top-10 and the training dataset of the target kernels during fine-tuning, since these are all labeled data points, and we calculate its speedup compared to the best design in the dataset. We run the offline evaluation five times and the online evaluation three times, then we report the mean results.
@@
The experiments demonstrate MoE’s effectiveness (shown in Table 1). We use the second design of the high-level gating network as it performs better. MAML could not perform well in this setting. It might be because we have many more source kernels compared to the paper that proposes to use MAML on this task (Bai et al. 2022). Different kernels might result in different directions of the meta gradient, leading to unstable training. ProgSG, which combines GNN and a language model, has a strong ability when the training data is sufficient, but it overfits in the data-scarce setting. During fine-tuning, the training loss is lower than 0.05, but the test loss is high. Comparatively, the MoE is more generalizable.
Hierarchical MoE is better than single MoE. In the online evaluation, on most kernels, the best of three single MoE models outperforms HARP, but different kernels favor different ones. For example, block is the best granularity for “Fd’,’ while graph is the best granularity for “Gemv,” “Ja,” and “Tr.” Comparatively, hierarchical MoE performs the best or close to the best on almost every kernel.
@@
Complex kernels. Since HLSyn kernels are generally simple, we test the model on 5 more complex kernels and verified the effectiveness of Hierarchical MoE. The 5 kernels include ”3d-rendering” and ”spam-filter” from the Rosetta benchmark (Zhou et al. 2018) and three self-constructed kernels that are introduced in the appendix. These are closer to real-world applications. Following previous settings, we train the regression model on HLSyn kernels and finetune it on 50 points per kernel on the new kernels. Table 3 shows the results. It is challenging for all models to perform well, but the hierarchical MoE model performs the best. It remains a future work to improve models on complex kernels.
@@
Table 4: Using MoE on various granularities in a single model. “N,” “B,” and “G” represent node, block, and graph.
To verify the effectiveness of our model design, we do ablation studies on different model structures. First, instead of aggregating the three low-level MoEs, we try applying MoE on the three granularities together in a single model, and we specify the detailed structure design in the appendix. Table 4 shows the results. When we use MoE on two or three granularities in a single model, the loss is not only higher than hierarchical MoE but even higher than the lowest loss when only using MoE on one granularity. Apart from selecting 50 representative data points by K-means for fine-tuning, we also experiment with random data split. Hierarchical MoE is the most stable model when it faces low-quality fine-tuning data in the random split. If we stack MoE on various granularities in the same model, the structure might be too complex and thus unstable to train on a small dataset. The results demonstrate the effectiveness of hierarchical MoE.
@@
Second, we want to examine whether the hierarchical MoE benefits from combining the three granularities. To answer it, we still use the hierarchical MoE structure, but we only use one low-level MoE’s granularity. Table 5 shows the results. It verifies that aggregating the three granularities performs the best. Using only the node MoE or graph MoE also performs well in the online evaluation, and this might be due to the increased number of experts. Nonetheless, aggregating the three granularities further improves the performance.